New Highlights

News Items

New 150 grating (2022-Sep-01)

A new 150/7500 is available for science.

Grism issues (2021-Nov-24)

Initial remedial work on the grism transport mechanism has been succesful. The grism is fully functional. We will continue to monitor this mechanism.

Grism issues (2021-Nov-1)

Deployment of the grism at certain angles has become problematic. The main issue is the transport mechanism of the grism. This mechanism transports the grism from its carrousel (STOWED position) to the beam (DEPLOYED position). Depending on the orientation of the instrument (rotation of the instrument and elevation of the telescope), the grism does not reach the DEPLOYED or STOWED positions.

The effects of this issue are significant because the grism has to be transported in and out of the beam during MIRA, Mask Acquisition, change of spectroscopic setup, and for imaging. Therefore, we suggest that the observers plan carefully their observations so that changes of the grism are minimized during the night.

We are working to mitigate this problem in several ways. We are working on adapting the SAT so that masks can be acquired with the red channel by using the gratings as a mirror (zeroth order as it is done with Deimos). Moreover, we are developing a plan to perform a mechanical study to fix this issue in early 2022. A major refurbishing work of LRIS is planned to take place in 2023 to permanenty fixed this and other LRIS issues.

150/7500 grating out of service (2021-Oct-12)

The 150/7500 grating has been taken out of service due to a crack near a corner. We are starting the process to replace it.

Polarimeter Status (2021-Sep-19)

The LRIS polarimeter is being repaired. We expect to get it back from the optics shop by the end of 2021. Its recommissioning would take place at the start of 2022 and it will enter into service shortly there after. We will post news about it as soon as we can.

New LRIS Red CCD (Mark IV) (2021-Aug-19)

A new monolithic 4k x 4k thick CCD has been installed to replace the old Mark 3 LRIS red side mosaic (late April 2021). This new CCD has a faster readout and a similar QE as the previous mosaic.

Binning by 2 in the spatial and spectral direction is possible. However, windowing is not possible at the moment but we expect to commission this mode in 2022.

The new CCD is thicker than the old mosaic, therefore we recommend maximum exposure times of 20 minutes to avoid excessive contamination by cosmic rays.

LRIS Red Detector Update (2020-Aug-02)

We continue to experiece problems with the LRIS red side detector.

The LRIS RED detector has 4 video amplifiers, called "vidamp1" to "vidamp4". When we use "LEFT ONLY" mode, the two CCDs are readout by vidamp1 and vidamp3. When we use "RIGHT ONLY" mode, the two CCDs are readout by vidamp2 and vidamp4. When we use both LEFT and RIGHT modes, all 4 vidamps are used. At the time of writing, vidamps 1, 2 and 3 are ok. Vidamp4 is not functional. The result is that LEFT mode is ok, RIGHT mode is missing one of the two CCDs (the one readout by vidamp4). If LEFT+RIGHT mode is used, 1/4 of the mosaic (1/2 of one CCD) is missing.

The situation is rather unstable and we occasionally see failures of other amplifiers.

Most of the times it is possible to operate using the left amplifier, which means that the only effect is an extended readout times.

Occasionally, we see an entire CCD becoming unresponsive. In these cases, the effect is much severe: MOS observations cover only half of the field of view and polarimetry is no longer possible. Please contact your Staff Astronomer for the latest updates ahead of your observing run.

LRIS Red Detector Update (2020-Feb-13)

We have recently been experiencing problems with the LRIS red side detector. While the details and history are a bit convoluted, the current situation is that we have one bad amplifier on the bottom chip (as shown in the ds9 live display at the instrument). The workaround for this is to use the single amp readout mode (with the right side amplifiers). This yields a complete image with no bad sections at the cost of about double the readout time. The readout time (regardless of windowing) is approximately 195 seconds for 1x1 binning and 60 seconds for 2x2 binning.

For observers who are not interested in data on the top half of the bottom chip (i.e. long slit observations of small or point source targets at the "slitb" pointing origin), you can use the dual amp readout mode and get faster readout times.

This problem had evolved with time. While the situation above appears stable for now, it may get worse. We will update observers if that happens.

DVD writer no longer available for data backups (2016-Jan-26)
The Keck Observatory is phasing out the auto DVD writer sometimes used by observers for data backups. Options to launch the DVD autobackup gui are no longer available from the background menu. Because the Keck Observatory Archive (KOA) now ingests all raw data from every instrument at the observatory, observers are encouraged to use KOA to retrieve a copy of the data following your observing sessions. Instructions and links to possible methods of data backup including KOA, scp, and rsync are found at: BackingUp.

New calibration spectrum for Fe-Ne lamp and for deuterium lamp (2015 Feb 13)
The calibration spectrum for the new Fe-Ne lamp and the corresponding line list is now available in our calibration page. Note that this is not the standard Fe-Ar lamp that most instrument use. Calibration spectra have been provided only for the 1200/3400 grism, as this is the only one that benefits from the increased blue coverage and the large number of lines. Two spectra are provided, for the blue and red side of the range covered by this grism.
Our spectroscopic flats page now provides a reference flat field spectrum for the new deuterium lamp, obtained with the 400/3400 grism, 1 arcsecond slit, and 30 seconds of exposure. This spectrum should only be used for wavelengths shorter than 4000 Angstroms, as it displays strong emission lines at longer wavelengths.

New calibration lamps (2014 August 21)
A number of new calibration lamps have been added to the LRIS calibration system: Two new red side arcs (Xe and Kr), and a new blue side flat field lamp (Deuterium, 2H). Calibration spectra for the new red lamps are available on our web pages. The new 2H lamp is especially designed to provide illumination below 4000 Angstrom. It produces significant illumination at redder wavelengths too, but there are many emission lines. Recommended exposure time with the 400/3400 grism is 40 seconds; with the 1200/3400 it is 300 seconds. The lamp takes about 3 minutes to warm up.

There is an additional blue arc lamp already installed (FeAr) but so far we have not been able to find a suitable comparison spectrum for it. We suspect that the lamp might be damaged and we don't recommend using it.

New command to change red side grating to reduce wavelength uncertainties (2014 June 9)
The grating tilt mechanism of the red side has recently shown evidence of non-repeatability. The position of the calibration lines has been observed to change by as much as 100 pixels upon removal and insertion of the grating. This has been tracked down to an incorrect calibration of the grating tilt encoder. To fix the problem, we encourage observers to use the "grating" script to change the grating. As an example, if you want to put the 1200/7500 grating in the beam, and set the wavelength to 8000 Angstroms, the syntax is:
 grating 1200/7500 8000 
If the wavelength is for a slitmask, rather than a longslit, the syntax is:
 grating -mswave 1200/7500 8000 
This command contains a forced homing of the grating tilt mechanism. Repeated tests showed that homing the stage before setting the wavelength seems to eliminate the problem.

DVD writer for data backups will be phased out (2014 March 7)
The Keck Observatory is phasing out the auto DVD writer sometimes used by observers for data backups. Although there is no time line for the removal of the DVD writer, we will no longer troubleshoot or maintain the DVD writer. Because the Keck Observatory Archive (KOA) now ingests all raw data from every instrument at the observatory, observers are encouraged to use KOA to retrieve a copy of the data following your observing sessions. Instructions and links to possible methods of data backup including KOA, scp, and rsync are found at: BackingUp.

New Gold Coated Gratings (2013 March 21)
To take full advantage of the red side detector upgrade, three gold coated gratings were released for observer use. These gratings include:

These gratings are listed on the Grating and Grisms page and are available in the pulldown menus on the LRIS configuration form.

New telescope and guiders focusing procedure (2012 Nov 8)
For some time now we have been collecting data on the performance of the IQM focusing procedure, which is about ready to be deployed. To reduce the remaining uncertainties we need to run this procedure on every LRIS night. As a consequence, the Observing Assistant will routinely run IQM at the beginning of the night, in twilight, and will coordinate with the observer on a window of opportunity to run IQM a second time during the night. The expected running time of the second instance is about 5 minutes. Note that this is run after the standard focusing steps are completed, namely a MIRA and/or an autofoc. Those steps would be required in any case.

SAT slitmask alignment tool updated (2012 Jan 14)
The astrometry used to predict guider fields based on target coordinates was recently updated and observers may again use the Slitmask Alignment Tool (SAT) to coarse align slitmask. Due to mechanical flexure, the uncertainty in predicting a guider field is 1.5 arcsec, and new hardware schedule to be installed in February should help minimize the flexure. Please see the SAT web pages to learn how to coarse align slitmasks using the SAT.

New Slitmask Names are (Temporarily) Being Assigned To Your Masks (2011 Nov 08)
For security reasons, LRIS mask names are now being overridden with automatically generated names to prevent the possibility of a malicious filename causing trouble with the database. Your support astronomer can change the slitmask names back to their original values on the day of your run; however, when completing the LRIS configuration form you must use the nonsensical names assigned by the slitmask ingestion system. We hope to have this issue resolved in the near future so that legal mask names assigned by the observing team will be accepted.

Keck starlists can now contain RA/Dec offsets (2011 Oct 24)
In response to a recent observer request, the MAGIQ guider software has been updated to allow observers to specify RA and Dec offsets to be applied after a target has been centered up. The intention is to permit starlists to contain entries for alignment stars with known offsets to a faint science target which may not be visible on the guider. As described in the starlist documentation, the new keywords to include are raoffset and decoffset. The offset values listed in the starlist file can be read directly into the MAGIQ offset fields, thus avoiding the need to verbally relay telescope offsets to the Observing Assistant (with the corresponding risk of a miscommmunication). If you have any questions about how to specify offsets in your startlist file, please contact your Support Astronomer for assistance.

IQM for focusing the telescope (2011 Oct 21)
The best focus determined using the MIRA software is good with LRIS on Keck 1, but we can do better. We are now employing IQM on LRIS to better focus the telescope during the night. We have demonstrated that IQM accurately corrects for piston moves on the secondary but not the tilts. The secondary tilts do not change over the course of the night, and so we have developed the following procedures to focus the telescope.

This is a transition phase to only using IQM during the night.

The reason why we believe this should provide a better focus for the scientific observations is that LRIS instrument focus is a function of elevation and rotation. Because MIRA relies on the science camera being in focus, MIRA will suggest different secondary piston positions for different locations on sky. In contrast, the guiders have a focus dependence on rotation only, and the magnitude is 10 times smaller than those measured with LRIS.

Grating Monitor (2011 Sep 22)
Because XLRIS sometimes does not report the correct grating, a new script called Grating Monitor was released and will launch with the rest of the software. We have found that both XLRIS and the xshow command employed by the LRIS Motor Status gui will report gratings in the same grating port that were installed the day before instead of the current grating. We have tracked the problem down to the keyword broadcast code and suspect a memory leak in the code. Until this issue is addressed, please check the Grating Monitor for the currently selected grating and restart XLRIS when needed.

Generate Mask Starlist (2011 Sep 18)
This is a new tool available from the pull down menu that queries the slitmask database for the mask RA, Dec, and PA specified in the design file, and then it generates a starlist for the masks. It is used to build a starlist for the slitmasks that are currently loaded in the instrument and is to be used when preparing for the night's observing in the afternoon. So far, the software has been tested on a handful of configurations with 100% success. By default the starlist is written to the numbered account home directory on the lrisserver. This new script is available on the pulldown menu under LRIS Utilities -> Generate Mask Starlist.

Guider Flexure update (2011 Sep 15)
Exhaustive flexure testing has revealed that the observed guider flexure is due to mechanical shifts in the guide camera mount. All other optical components in the system are stable. Mechanical engineers are developing a solution to stiffen the camera mount.

New AUTOSLIT version released. (2011 July 26)
AUTOSLIT 3.20 was released by Judy Cohen and Patric Shopbell. The new version depreciated the use of pickoff mirrors, plots both the red and blue side chip gaps, added slit overlap checks for blue and red gaps, and outputs a ds9 regions file that may be overlayed on an image. The new version will operate on Mac OS 10.6 Snow Leopard, RedHat Enterprise Linux with gcc 3.x and 4.x, and Oracle/Sun Solaris 10 with SUN cc. Please see the AUTOSLIT web doc to download the new version.

Remember to Focus LRIS Before Running MIRA (2011 July 07)
When the OA runs the MIRA script to focus the telescope, a key assumption is that the instrument is properly focused; it it's not, then the derived telescope focus may be wrong. As the observer, it's your job to ensure that the blue camera is correctly set to the G-band imaging focus before the OA runs MIRA. This also means that you need to remember to measure the direct imaging G-band focus in the afternoon, even if your observing program will not require any imaging observations! Please consult your Support Astronomer if you have any questions about focusing LRIS.

Use the Right Dichroic When Imaging Slitmasks (2011 July 04)
Observers should be aware that changing the dichroic will shift the image on the blue mosaic. When acquiring afternoon direct images of your slitmasks to define the alignment box locations, you should be sure to use the same dichroic that you will use on-sky for slitmask alignment. Generally, this can be the same dichroic you will use on the blue side for spectroscopy. Please note that the Slitmask Imaging Tool has controls which allow you to select the dichroic when acquiring your afternoon slitmask images.

LRIS Offset Guider Flexure (2011 June)
During the last LRIS run, one of our observers found that the Slitmask Alignment Tool was not correctly predicting the guider field at the coarse alignment step in the procedure. At first, it appeared that the guider field center was not well defined, but the mismatch was different at several rotator values. We have since discovered that the offset guider has developed flexure that is large enough that it impacts coarse aligning slitmasks using the SAT. Over the past week, we eliminated various sources of flexure in the lris offset guider system, but we have not found the root cause.

Because of the increased flexure, the support staff is sadly recommending that slitmask coarse alignments be completed using older methods instead of using the coarse alignment tab on the SAT. Please see these instructions for a description of how to coarse align a mask. Long slit observing is not impacted by this fault. Multi-slit observers need to be aware of this issue and it is recommended that you check the mask alignment when the rotator is changing rapidly (once and hour).

MIRA vs IQM (2011 April 14)
Until July, the OAs will be acquiring both IQM and MIRA data when focusing the telescope. IQM is a second method of focusing the telescope and will eventually replace the need for MIRAs (the standard focusing method) during the night. In the short term, this will add a couple minutes to telescope focusing, but in the long term, IQM will save five minutes per telescope focus. If you have any questions, please contact your support astronomer.

Red Detector linearity improved (2011 April 14)
The LRIS Red side detector linearity was improved such that all four amplifiers are linear to about 60,000 ADUs. Recent changes to how the array was clocked improved the linearity especially for detector 19-2 (top half of mosaic) where the typical long-slit spectrum is placed. Many thanks to Bob Kibrick for testing and further improving LRIS Red.

New Slitmask Alignment Tool released for use (2011-Jan-01)
To celebrate the new year, we released a new Slitmask Alignment Tool (SAT) for observer use. The new SAT are more efficient at aligning masks and has several advantages over the old scripts and IRAF/xbox routines. First, the SAT displays fits to all objects and boxes simultaneously along with the residuals plot and calculated offsets. This helps observers more easily identify stars that should be excluded and included in the fitting process. Second, predicted guider images may be used to coarse align a mask (DEIMOS style), saving image readouts and time spent identifying alignment box stars in LRIS images. Third, the IDL based routines simply execute quicker than the IRAF scripts.

With the release of the SAT, we have updated the slitmask alignment procedures. To review the new routine please see: Slitmask Alginment Procedure

Below are screen shots of the tool for the three primary observing uses.

LRIS Red Mark II commissioned (2010-Dec-03)
The LRIS Red Mark II dewar and detector upgrade achieved first light, performed well, and is now ready for regular nighttime use. In comparison to the previous system, the new detectors are cosmetically superior and have fewer charge traps. Readout times are relatively longer with readout times of 127s and 58s for 1x1 and 2x2 binning, respectively. The full field of view is available with the new system and slitmask should be designed to take advantage of the entire FOV for LRIS Red. The performance of the new detectors is not yet completely characterized, but will be posted by January 2011. Contact your support astronomer if you have questions.
A new dewar and detector system will be installed

LRIS Red III - status (2010-Nov-22)
A new dewar and detector system will be installed starting 29 November. This new system replaces the failing LRIS red detectors experiencing a CTE problem. The new system is scheduled for first light on 2 December and observers from December on will be able to use the new system. The full field of view is available with the new system and slitmask should be designed to take advantage of the entire FOV for LRIS Red. The performance of the new detectors is not yet characterized, but will be posted by January 2011. Contact your support astronomer if you have questions.

CTE now evident in LRIS Red CCD Mosaic Last Amplifier (2010-Oct-19)
After examining the cosmic ray detections in 15 min darks, we have determined that about ~5% of the flux is smeared in the amplifier #4 (lonslit position) vs 25% in amplifier 3. Visually, the smearing in amplifier 3 extend 50 pixels, while in amp 4, the smearing appears to extend around 20 pix from the peak flux of a cosmic ray detection. Please see for an example spectrum of a standard star.

CTE now evident in LRIS Red CCD Mosaic Last Amplifier (2010-Oct-18)
LRIS Red data now exhibit the CTE problem in the last “good” amplifier in the LRIS Red mosaic. The location of the last good amplifier is the portion of the CCD mosaic where longslit observations are acquired. Thus, every scientific program with LRIS Red is now impacted by the CTE problem. For a review of the CTE problem, please see past LRIS News items.

A sample 900s dark is available upon request. The cosmic ray detections may be used to help you assess how severely the CTE problem impacts your science program. The CTE problems make faint object extraction and sky subtraction more difficult, but may still be possible. It may be possible to avoid using the red side, if with the right dichroic and grism, your science goals may be achieved using the blue side. Please let me know if there are any blue side calibrations which may help you assess the possibility of using the blue side.

For now, our plan is to continue observing with the system as is. Observers should use LRIS in one amp or dual amp mode depending on whether their observational program is longslit or multislit, because at the very least, these modes can suppress the CTE problem. If you are unable to use LRIS with the current condition, please contact your support astronomer.

LRIS Blue shutter repaired but FOV shifted (2010-Jul-1)
We have replaced the failed blue shutter with one of two spares, and are now in the process of repairing the failed shutter. The replacement shutter operates normally.

The process of replacing the shutter includes removal of the dewar and camera barrel optics assembly. After re-installing the optics and dewar, we re-aligned the dewar. The rotation of the dewar is nearly identical, but there is a 25 pixel shift (3.4 arcsec) in detector X. Because of the shift, the traditional longslit position will be closer to the gap on the blue side only. Slitmasks that were designed with slit at the edge of the gap will be miss 3.4 arcsec of the slit.

For the current LRIS run, observers need to take into account the spatial shift. Following the current lris run we will attempt to adjust the dichroics shift the FOV to the previously "good" position.

LRIS Begins Full-Time VNC Operation (2010-Jun-27)
Starting with the July LRIS run, we ask all observers to start up their LRIS control software within VNC desktops. (VNC: ) is a technology that allows windows to be shared with observers at other computers and has been used successfully for years at Keck to provide mainland observers with access to Keck instrument controls. By expanding the use of VNC to all observers, we will thus have a single startup procedure for observers to follow regardless of whether or not they will operate with mainland observers. Operating under VNC also makes it easier for support staff to assist observers with troubleshooting and allows our observing assistants to monitor the status of observations during the night. Your support astronomer will be happy to assist you with starting your LRIS software under VNC.

Blue Side Shutter Failure (2010-June-16)
The Blue Side shutter has failed, and will not be functioning for the duration of the current LRIS run. The shutter was forced open and observers have options to observe in a shutter-less mode or to use the trapdoor as a shutter. Please review the strategies for failed shutter observing. As soon as LRIS comes off sky we will replace the shutter with a spare, and have LRIS repaired before the July observing session. To replace the shutter the dewar must be removed from the instrument to gain access, thus we are waiting until the end of the current observing run to repair it. When the shutter is stuck open, the data look streaked in the readout direction as seen in this direct image of a slitmask.

Compass Rose - Welcome Back (2010-May-20)
At long last, the compass rose returns for LRIS. Because the image orientations are the same on the red and blue sides, only one compass rose is needed. The updated compass rose is displayed on the blue side display monitor to assist with field identification and slitmask alignments.

readmhdufits IDL script bug fix (2010-May-20)
Found a bug in the IDL code readmhdufits used to combine the individual amplifiers into a single FITS image. We found that DETSEC and DATASEC were indexed differently with DETSEC starting at 1,1 and DATASEC starting at 0,0. This resulted in shifts of 1 pixel. The bug is now fixed. Please download the latest version from the readmhdufits web page.

Addressed: Objects drifting out of slits (2010-May-5)
Last week, we completed some repairs of the LRIS offset guider nikon camera lens to address a drifting problem with the LRIS offset guider (see news item below). Through flexure testing, we found that the camera lens and mount were loose and possibly damaged. Even though we have now securely attached the fmount to the camera body there is still a small amount of movement. The motion we observed before our fix was a "flop" with rotator angle which would result in a rapid 1.5 arcsec "drift" on the science/guider fields (enough to drag science objects out of the slits). AT this point, the "drift" problem is significantly reduced, but not eliminated. Observers should remind their OA about the problem and ask the OA to try using the slit guider if a "drift" in guiding is observed. Two possible sources of flexure persist in the offset guider. First, the camera lens is still loose as noted above. Second, we found that one of the fold mirrors in the optical path will flex on the order of an arcsec with the instrument at horizon. This is a more minor source of flexure and has been present since the new LRIS MAGIQ offset guider was commissioned. Both flexure sources will be addressed the next time LRIS comes off sky.

Objects drifting out of slits (2010-April-16)
During the most recent LRIS observing run, observers and observing assistants discovered that objects were drifting out of their slits. When guiding on the offset guider, we have found that objects could move by as much as one arcsec in an hour on the LRIS science camera while the guide star position is unchanged. We have not observed drifts when using the slit guider for guiding. We suspect new flexure in the offset guide camera is contributing to the observed drifts. We are advising two courses of action to correct for the drifts. First, you can check your alignment every half hour. Removing the grism from the optical path and re-imaging the field will result in a few minutes of lost time, but the checks will update the science object position. A second option is to use the slit guider to centroid on a bright star seen on your slitmasks. Bright stars may not be available for guiding on slitmasks, so this may not work for every science field. The procedure for the second option is to slew to your field; course align the mask using the offset guider; have the OA to start guiding using the slit guider if possible, and then continue with the course alignment. Keck staff continues to monitor the problem, and will investigate the guider hardware when LRIS comes off sky at the end of April.

Calibration lamp location moved (2010-March-16)
The calibration lamps use to be mounted on the movable guider stage. Recently, the movable guider was removed after making a permanent switch to using the new MAGIQ offset guider. The calibration lamps are now mounted on either side of the LRIS FOV. Because of the new mounting scheme, lamp intensity is different than past observations. Please carefully check that the lamp intensity in your calibration data is appropriate for you science needs.

LRIS Red Warming (2010-March-16)
We have addressed one problem with the LRIS Red dewar, and we are now cooling it down for use. We have leak checked the system and used an RGA during cool down to determine where the leak is but as of yet we have not found it. We have replaced some o-rings which may have been at fault but leak checks before and after the o-ring replacement did not indicate an improvement. There is a chance that we will again encounter the leak, which will mean that we will again have to warm and pump the dewar before getting back on sky. Observers should keep in mind back up projects that may be completed with the blue side only.

LRIS Red Warming (2010-March-12)
Over the last three weeks, LRIS Red has unexpectedly warmed twice. The warm-ups are accompanied by an increase in vacuum pressure, and we suspect that we have a small vacuum leak. With the dewar at ambient temperatures, we have leak checked twice, but we have not observed a leak. The last warm up occurred on 9 March, and following the warm up, we leak checked the dewar and o-rings on the pressure relief valve were replaced. No leaks were discovered, and we suspect that the leak is only present when cold. The dewar was pumped down for a few hours and then cooled in time for Wednesday nights observing. The following day we topped off the LN2 three times before starting Thursday night's observing session, but the cryo lasted only until midnight when we noticed that the dewar temp had increased to -120 C. We suspect at this time that either a vacuum leak is reducing the hold time or that we did not pump on the dewar long enough during the last unexpected warm up. Our plan following Thursday night's observing is to allow LRIS Red to warm to ambient temps and pump on the dewar for a few days to ensure an adequate vacuum. This impacts your observing session because it means that LRIS Red will not be available. Please plan accordingly.

LRIS Calibration Widget Released (2010-Feb-18)
The new LRIS Calibration Tool automates the process of acquiring calibration data (arcs and internal flats) for a series of longslits or slitmasks. See these instructions on its use.

LRIS Red CCD Mosaic Stabilized. (2010-Feb-18)
To stabilize the LRIS Red Detector problem, WMKO and UCO/Lick staff have added new electronics to LRIS to maintain power during instrument change outs and new software that will supply a steady heater current. This ensures that the CCD temperature remains at or above the operating temperature during instrument swaps. With these additions in place, we have stabilized the operational state of the LRIS red CCDs; however, this will not recover the three degraded amplifiers, which remain unusable for science.

To take advantage of the one remaining good amplifier, observers acquiring data in longslit mode should select dual-amp readout mode to save readout time. Observers acquiring imaging or multislit data should select single-amp (right) readout mode with the right-side amplifiers to obtain good images over 1/2 of the mosaic at the expense of doubling the readout time to three minutes per exposure (unbinned). Several observers have questioned whether it is possible to obtain anything on the "bad" CCD in one-amp mode. The performance of the good CCD on the good amp (1-13 R) is unchanged since commissioning: linear to 64K ADUs, gain=0.916, and RN=3.10. In contrast, the performance of CCD 1-12 R will be non-linear, gain=1.3, and an RN > 17 e-. Observers may use the "bad" side at their own risk.

LRIS Red CCD Mosaic Loses Another Amplifier (2010-Jan-05)
Images acquired with the LRIS red mosaic today show that a third amplifier (out of four) has now gone "bad", as shown by trailed cosmic rays. The only remaining "good" amplifier is the one on which the longslit spectrum falls in dual-amplifier readout mode. We suspect that thermal cycling that occurs when LRIS moves on or off the telescope is responsible for the worsening problem with the detector mosaic and will take steps to minimize such cycling in the future. For now, we make the following recommendations to observers:
  1. In longslit mode, use dual-amp readout mode to save readout time.
  2. In imaging or multislit mode, use single-amp readout mode with the right amplifiers to obtain good images over 1/2 of the mosaic at the expense of doubling the readout time to 3 minutes per exposure.
Please note the following:

IQM testing (2009-Dec-14)
With the deployment of MAGIQ hardware on LRIS, we now would like to begin acquiring MAGIQ Image Quality Measurements to test whether we may be able to correct for changes in telescope focus using MAQIQ IQM instead of using MIRA during the night. To make the necessary tests, we are asking observers to allow the OAs to acquire MAGIQ IQM data before and after each MIRA during the night. The time required to take the IQM data will be less than less than five minutes in addition to the amount of time needed to run the standard MIRA telescope focus. Please consider this request for your upcoming science runs. For further information, please see IQM vs. MIRA Testing

LRIS Red CCD Image Readout Problem Update (2009-Dec-10)

Recently, the performance of 1-12 has degraded further. The LRIS Red team at UCSC has tried modifying 1-12 voltages and timing schemes to again recover it for scientific use, but so far, the CCD is not useable for science observations. For the upcoming runs in December and January, half of the LRIS Red mosaic will not be useable for science observations, and observers should plan accordingly. Imaging or multislit spectroscopic science programs requiring the entire FOV are impacted. If you need to redesign slitmasks for the upcoming run, please submit them soon. Please see "LRIS Red CCD Image Readout Problem Update (2009-Oct-30)" in the news below that has example output from the Autoslit slitmask design software that shows where to safely place slits for red side observations. As a reminder, the region on which the longslit spectrum typically falls is not affected by the CTE problem as longslit spectra typically fall on detector 1-13 (functioning normally). Red side polarimetry programs will be severly impacted as one of the beams falls on the non-useable side.

LRIS Red CCD Image Readout Problem Update (2009-Nov-9)

The LRIS Red CCD and electronics were shipped back to the W.M. Keck Observatory 4 Nov. 2009 and was reinstalled in LRIS in time for the upcoming LRIS run starting 9 Nov. 2009. Although the charge transfer efficiency problem is not fixed, the UCSC team has made tremendous strides to recover a half of the detector that is closest to the gap. This means that in two-amp readout mode about 2/3 of the illuminated area of detector 1-12 is usable. As a reminder, the region on which the longslit spectrum typically falls is not affected as it falls on detector 1-13 (functioning normally). Imaging or multislit spectroscopic science programs requiring the entire FOV are impacted by this problem. For observers who want or need the full LRIS FOV to achieve their science goals, the array may be read out in single-amp mode to recover the entire LRIS Red FOV at the cost of increasing the readout time, as indicated in the table below.. Although the UCSC team was able to recover half of the CCD, the serial CTE, gain, effective readnoise, and linearity is degraded relative to values determined at commissioning. Below are preliminary performance results which will be confirmed after the system is re-installed in LRIS. These performance numbers mostly pertain to the 1-12A half of the CCD. The performance of the second half (1-12B) has not been quantified completely but is in general worse. If observers want to use the full array, we are recommending that detectors be readout in single-amp mode. In single-amp mode, the 1-12A specifications will apply to both A and B sides of detector 1-12.

The gain and readnoise for 1-12A are now 1.224 (e-/DN) and 6.52 (e-/rms), while at commissioning three months prior the values were 1.022 (e-/DN) and 4.67 (e-/rms). The UCSC team also measured gain and readnoise for 1-12B: current - 1.171 (e-/DN) and 12.9 (e-/rms) vs before 0.955 (e-/DN) and 3.79 (e-/rms).

The CTE for 1-12A as measured using deferred charge (ratio of signal in the last column to that left behind in the over scan) is not constant, but it depends on signal level. The main effect of the smearing will spread 13%, 2%, and 6% of the object flux in the spatial direction for signal levels of < 250, 1000-15000, and >15000 DN, respectively. This flux will be smeared into the sky where it will be difficult or impossible to recover. The bad CTE at high intensity levels will be a problem for flats as there are many sharp features in the flats and those will not be recovered well with bright flat exposures that smear.

Detector 1-12A is linear to 50,000 DN, while 1-12B appears to be non-linear. Once LRIS is installed in the instrumen, we will be able to better quantify linearity.

AMPMODE   XBIN   YBIN	Readout time   Total (Erase + Readout)
--------  ----   ----   ------------   -----------------------
DUAL:L+R    1      1           83s       100s
SINGLE:L    1      1          153s       170s

DUAL:L+R    2      1           51s        68s
SINGLE:L    2      1           90s       107s

DUAL:L+R    1      2           45s        62s
SINGLE:L    1      2           81s        98s

DUAL:L+R    2      2           30s	  47s
SINGLE:L    2      2           49s        66s

LRIS Red CCD Image Readout Problem Update (2009-Oct-30)
The UCSC team checked the internal connections and they all appear to be good and replaced a cable. After cooling down and testing, the problem persists. In the short term, we have decide that the dewar will be shipped back to Keck Nov 4 so that the good ccd (1-13) may be used for the upcoming science run starting Nov 9. While LRIS is on sky, the UCSC team will have two parallel troubleshooting activities. Test an engineering grade detector in their test setup to determine whether broken connections can reproduce the problem. Second, they will play with waveforms on CCD 1-12 (poorly performing detector) while LRIS is installed. If the waveforms can be adjusted to obtain normal readouts, then 1-12 can be made available for use during the November run.

Observers who will be using slitmasks, should be aware which region is available on the red side. This slitmask image shows where slits may be placed to land on the good CCD on the red side. The image is example Autoslit3 output. You are not limited to using just the top half of a slitmask but you should consider the locations of your highest priority objects.

LRIS Red CCD Image Readout Problem Update (2009-Oct-26)
The dewar and associated electronics arrived at UCSC Monday morning. Upon arrival, the dewar was opened after backfilling with ultra-dry N2, and internal cabling and connectors were inspected. UCSC staff also examined the wire bonds on the CCD, and a series of photographs that show that all bond wires appear to be intact. Thus far, nothing obvious was discovered in the initial internal component investigation.

Late Monday, UCSC staff re-assembled everything, replacing a flex cable with a spare. The dewar is now on the vacuum pump, and the plan is to cool the dewar Tuesday morning with testing to follow in the afternoon.

LRIS Red CCD Image Readout Problem Update (2009-Oct-24)
The LRIS Red Dewar and electronics were removed from LRIS following the last night on sky in October and was shipped back to UCSC for further investigation by the instrument team. Upon arrival at UCSC, the dewar will immediately be opened and internal components investigated. In parallel, the electronics will be tested with their test dewar to ensure proper operation. At this time, we fully expect the dewar to return to Keck for use during the November run with at least one normally functioning CCD.

LRIS Red CCD Image Readout Problem Update (2009-Oct-10)
As most of you are aware, one of the CCDs in the new LRIS Red detector mosaic is experiencing a charge transfer efficiency problem. Troubleshooting has determined that the problem is internal to the dewar. Friday, we decided to send the LRIS Red dewar to UCSC to attempt a repair of the poorly performing CCD immediately following the current LRIS run ending the morning of Oct 24. The dewar and associated electronics will be removed from LRIS on Oct 24 and then shipped to UCSC on Oct 25. At UCSC, the dewar will be opened and internal components inspected. Small repairs will be completed if the identified issues are obvious and easily fixed. Repairs that will require parts with long lead times or otherwise deemed longer term will be deferred to a later date. If repairs are deferred, the dewar will be returned as-is to the observatory so that at least one working Red CCD is available for upcoming science runs.

At this time, we fully expect the LRIS Red CCD mosaic to be ready in time for observing during the November and December observing Runs. We expect the Red side to function with at least the currently 1 good CCD, and hopefully, a repair of the poorly performing CCDs will be successful. As with any repair mission, there is a small risk that the system will not be available during your observing session due to circumstances as yet unknown, but we firmly believe the risk of LRIS Red being unavailable is very low. We will of course notify observers of any changes critical to LRIS red observing, and please view the LRIS News pages for regular updates.

LRIS Red CCD Image Readout Problem (2009-Sep-14)
One of the two CCDs in the LRIS red mosaic is reading out incorrectly for undetermined reasons. The problem affects half of the red mosaic and is characterized by a general smearing of the image in the vertical direction, as shown in the bottom half of the accompanying image of a longslit spectrum. Data from the affected detector are probably unusable. Fortunately for longslit observers, the region on which the longslit spectrum typically falls is not affected. Our investigation so far has eliminated all of the easily-accessible components of the CCD readout chain as culprits. Further testing of dewar electronics will take place after the completion of the September LRIS run.

Updated Versions of multi-HDU Readers Handle Binned Images (2009-Jun-19)
The original versions we provided of the IRAF and IDL routines for reading the new LRIS images were unable to read data acquired in binned formats. We have now given both tools the ability to read binned data. Download the latest versions of the tools in the links shown below.

IDL Routine Reads multi-HDU LRIS Images (2009-Jun-17)
As a service to the LRIS community, we provide an IDL routine which will read mutli-HDU LRIS images and assemble them into a 2-D array, with optional bias subtraction. Download the routine here.

Sample LRIS Images Available (2009-Jun-17)
As part of the May 2009 upgrade to the LRIS Red Dewar, the instrument datataking system has been revamped to output data in a different format. LRIS images now consist of multiple "header data units" (HDUs), one per CCD output amplifier. These images require different handling from the old IRAF images, which were written as simple FITS files combining data from all amplifiers. As a service to the LRIS user community, WMKO is providing a set of sample images to allow observers to work with the new data format. The collection can be downloaded here and includes both raw, multi-HDU images as delivered by the datataking system and simple FITS images generated using an IRAF task.

KECK.LRIS IRAF Package Updated for Red Upgrade (2009-Jun-09)
The keck.lris IRAF package has undergone a complete revision to support the new image formats now generated by the LRIS red and blue datataking systems. See this article for full details. Users may download the entire keck package here.

LRIS Red Upgrade: Big Red Arrives (2009-May-14)
The LRIS Red Upgrade project started the installation phase earlier this month. In the last two weeks, the old LRIS red camera and associated electronics were removed from LRIS. The new camera and dewar arrived today and is headed to the summit on Friday. Thus far, the installation is on schedule with first light on slated for June 6 & 7, and the first official science night scheduled for June 14. Observers have been asking about the improved FOV and spectral coverage. At this time we have not verified the pixel scale or increase in coverage. The predictions are that the pixel scale will be around 0.13 arcsec/pix with the change from 24 to 15 micron size pixels. Given the mosaic detector size, the new camera should increase the spectral coverage by around 25%.

Blue side video board noise (2009-March-30)
The video processor board on the left most CCD of LRIS Blue failed and was replaced with a spare. The spare board exhibits low level noise as seen in this bias image. The stddev in the bias is 2.7 in the noisy amp relative to 2.1 for the right most amplifiers. We are in the process of repairing or replacing the old video board. The failed board was sent to the manufacturer for assessment.

Red Side Halos Problem fixed (2009-March-20)
We removed the LRIS red camera and found some condensation on both sides of the dewar window. The dewar window was cleaned, and the dewar was pumped out. Following cool down, the halos reported on 2009-Feb-15 were no-longer detected.

Starlists Now Allow PA Info (2009-Feb-20)
For years, observers have been forced to provide their position angle information to the OA verbally. The new starlist file format now permits you to enter the position angle for your target using the ROTMODE and ROTDEST keywords. Here is an example of how to specify the position angle:
# slitmask with specified rotator position
PA_targ_name    00 24 00.00 +16 00 00.00 2000.0 rotdest=123.45 rotmode=pa

# target observed at parallactic angle
VA_targ_name    12 34 56.78 +01 23 45.67 2000.0 rotdest=0.0 rotmode=vertical

Red Side Halos (2009-Feb-15) -Fixed 2009 March 20
Images of point sources on the red side only have halos at a level of 1% of the source peak flux. An image of the grid of holes mask shows clearly this effect. This effect is also present in spectra on the red side. We suspect at this time that condensation is present on the red side dewar window. The Red side camera will be removed in early March to address this problem.

New Slit Guider Camera commissioned. (2009-Feb-15)
A new slit guiding camera was commissioned for use during engineering nights on 12 and 14 Feb. The new slit guider is rotated such that the slit is 29.5 degrees relative to the columns on the detector. The new camera is sensitive enough to reach a signal-to-noise of 10 in 1s on a star of v-mag=19.5. The unvignetted FOV is expected to be 1.6 arcmin diameter, and at 40% vignetting, the FOV is expected to be 2.2 arcmin. For observers who use pickoff mirrors, the new default pixel location on the guide camera for a pickoff mirror guide star is 459 618. A new slitb pointing origin was also defined to move the star away from a bad column on the red side detector. An image of M13 is available here. The dark line going through the field is a 1.0 arcsec wide long slit.

New Hatch Commissioned but has light leaks (2009-Feb-15)
All observers are advised to acquire internal calibrations after the dome lights are turned off because the hatch has light leaks. As part of the new guider upgrade, a new hatch was needed to house all the new guider optics. The new hatch was tested during engineering nights on 12 and 14 Feb, and it was determined that it has light leaks. An internal G filter pseudo-flat acquired with all internal lamps off, hatch closed, and dome lights on revealed the light leaks seen in the image here. The peak flux from the light leaks is ~3000 ADU, and it varies as a function of rotator position, telescope elevation, and telescope azimuth. This problem will be address in either early or late March.

831/8200 Grating available (2009-Jan-22)
We received and commissioned a new 831/8200 grating. The grating is available for general use, and there is an option to select the grating in the LRIS configuration form.

LRIS slit guider upgraded (2008-Dec-23)
We have commissioned a new slit guide camera and electronics for LRIS. This was an emergency commissioning of a new guider camera because the old slit guider camera was flooded with glycol after a coolant connector seal failed. The old camera was a Photometrics CH250. The new camera is a Photometrics CH350. New electronics are part of the commissioning and the electronics are not inside the electronics bay. The new camera has a wider FOV with significant vignetting in the corners (see image). It is also a more sensitive camera with a higher gain and less readnoise than the old camera by a factors of 3 and 4, respectively. With the old guider, objects were centered at pixel 191, 143. Last night's commissioning result show that objects should now be centered at pixel 277, 290. In the MAGIQ software, Observing Assistants should see two cameras: one for LRIS lris50m (as before), and another called lrisSlit for the slit-guider. Both MAGIQ and XGUIDE were updated and released for use.

831/8200 Grating status (2008-Nov-13)
A new 831/8200 grating was ordered and is scheduled to arrive in January 2009. It was determined that the old grating could not be used inside LRIS due to the risk of additional fracturing. The new grating will be made from the same master grating as the original LRIS 831/8200 grating.

Desktop Menu Returns (2008-Nov-13)
The desktop menu or background menu is now used to run the instrument, replacing the LRIS workspace menu that appeared in the upper left corner of every display.

New IRAF tasks make LRIS mask alignment more convenient (2008-Nov-03)
Three new tasks are now available in the keck.lris package to assist observers with aligning masks: All three tasks presume that the file listing the box coordinates is named, where slitmask is the name of the slitmask as shown in the SLITNAME keyword.

Grating 831/8200 damaged and unavailable (2008-Oct-9)
The 831/8200 grating is damaged and has been temporarily taken out of service. Three fragments, each roughly one cubic inch in size, were found in the corners of the grating. We are in the process of assessing the grating substrate to determine the risk of future fragmentation, and we are exploring options for repairing and replacing the grating. If we determine that the only solution is to replace the grating, the time frame for replacing it is 10 weeks. If we determine that the damage will not impact science and there is no risk of future fragmentation, then the time frame for repair is 6 weeks.

Web page makeover (2008-Aug-19)
All LRIS web documentation was updated to the new instrument page format. If you find content or pages missing, please contact your support astronomer.

New I band filter installed (2008-Jul-11)
A new I band filter is available for use with LRIS. This filter is an interference type filter with a relatively square passband with FWHM from 6986-8204 Angstroms. Please see the filters on-line documentation to view the transmission curve.

NB4000 filter has a long-wavelength leak (2008-Jul-10)
Observers noted that the NB4000 filter has developed a long-wavelength leak past roughly 7000 Angstroms. Engineering data of this filter was acquired on 8 July 2008 to characterize the long wavelength leak. Spectroscopic flats of the internal halogen lamp using the 1 arcsecond long-slit and 150/7500 grating tuned to 7000 angstroms were acquired with and without the filter in beam. The filter spectrum was divided by the flat to produce a relative transmission spectrum. Please see filters to view the plot.

New MAGIQ Guider Software released (2008-Jul-7)
We are in the process of upgrading the acquisition and guide cameras on all WMKO instruments, and LRIS has been selected as the next instrument to receive a hardware upgrade. An important element of our upgrade strategy is to first deploy the MAGIQ software for operation with the two existing LRIS guide cameras. The new software provides all of the functionality of the existing acquisition and guiding software while improving routine tasks such as star list management, field identification, acquisition, and guiding. Successful testing of MAGIQ software with the existing slit viewing and offset guiders was completed on May 27 and June 27, respectively. The MAGIQ software is released for use, and you are encouraged to use MAGIQ during your observing sessions. The hardware portion of the guider upgrade will take place at a later date.

Slitmask Design Software (2008-Apr-18)
At this time, it is recommended that observers use the Autoslit software to design slitmasks for LRIS. Dr. Arjun Dey and his observing team compared the mask designs that were output from both autoslit and the ucsc lris package and found significant difference in the slit locations. The specific recommendation by the author of the UCSCLRIS package is that observers use the AUTOSLIT field mapping until the UCSCLRIS software is updated.

FVWM with LRIS (2007-Aug-29)
FVWM is a multiple virtual desktop window manager for the X windows system which is now in use with all LRIS accounts. With the switch from open-windows to FVWM, a couple of other changes were required. Most instrument operations are now found on the "LRIS workspace menu" which is automatically launched on startup and will be located in the upper left corner of each display. The "LRIS workspace menu" replaces the open-windows background menu, and is similar to the "DEIMOS workspace menu." Second, clicking on the virtual desktop will bring up a FVWM background menu which is used mostly to start VNC, start local xterms, start browsers, and instrument menus. The FVWM is not used to run the LRIS software, and there are no options provided on that menu to run the instrument. The web documentation has been updated with these changes, but please let your SA know of any inconsistencies.

Blue Side compass rose (2007-Aug-29)
A compass rose that works with the blue side figdisp display is released for observer use.

Grating offsets applied (2007-Aug-08)

Grating offsets are now in use so that the actual central wavelength is closer to the desired central wavelength. Each grating and port combination requires a different offset. All gratings in all ports have been mapped. For most grating and port combinations, the accuracy to which the grating is positioned is better than 10 angstroms. Gratings in port 3 may exhibit larger offsets of up to 50 A. In the past, observers' noted offsets from 100 to 1000 Angstroms. If you notice a large offsets of > 20 Angstroms in the desired central wavelength compared to the actual central wavelength, please inform you support astronomer and add a comment to your post observing comments form.

I filter broke. (2007-July-20)

On 18 July, the I filter was found broken inside the instrument. One of the largest pieces was lodged in the grism mechanism and was preventing movement of the grism mechanism. There is a 3-4 mm gouge in the back surface of the 600 line grism. Fortunately, the gouge is located at the very edge of the grism and this should not alter the grisms optical performance. Because the filter pieces tumbled inside LRIS for at least half a night, we are taking the time to inspect all optical surfaces that may have been exposed to filter fragments. In the long term, we expect to have LRIS ready for operation by the start of the next LRIS observing session which begins 8 Aug. LRIS will be without an I band filter on the red side until a suitable replacement is acquired.

ADC and LRIS together at last (2007-July-09)

Please see the ADC docs.

LRIS-ADC MODE: The default observing mode for LRIS as of semester 2007B is to use LRIS with the new Keck I Atmospheric Dispersion Compensator (ADC), regardless of what you requested on your observing proposal. The presence of the ADC changes the focal plane scale of LRIS by a small but measureable amount, meaning that:

(a) old slitmasks designed for use without the ADC should not be used with the ADC, and

(b) new slitmasks must be designed with software that accounts for the ADC.

If you plan to use LRIS slitmasks, then you must download the new, ADC-compatible version of Judy Cohen's AUTOSLIT software (v3.09) or the similar FABMASK package by Drew Phillips. Please contact your support astronomer for ADC-related questions.

LRIS Blue side shutter fixed, and CCD readout problem fixed (2007-May-6)
The blue side shutter was fixed between 31 April and May 5. A bushing on the shutter was worn. and galling in the bushing was causing the shutter to stick. A dry lube was applied to the bushing.

The CCD can be read-out using all four amplifiers following a repair of the CCD signal cable. The ccd readout problem experienced was that the second amplifier did not see the signal. This meant that the detector had to be read-out in single amp mode, which is slower. During the shutter repair, technicians discovered that the CCD cable had a busted wire. Re-soldering the wire to the connector has fixed the readout problem.

LRIS Blue side CCD readout problem (2007-April-18)
As of April 17, one of the amplifiers on the "left" CCD of the blue-side mosaic is failing to read out properly. Troubleshooting suggests that the problem is a bad cable between the dewar and the saddlebag. Since we have no spare for the cable, we are unable to repair this fault until the run is over. In the meantime, the workaround is to read out both blue CCDs using the "right" amplifiers only. The command to put the blue side into this mode is:
mb amplist=2,4,0,0
Readout times will be double the usual for two-amp mode on the blue side.

LRIS Blue side shutter is stuck open (2007-April-16)
The LRIS blue channel shutter has failed again, and work to replace the shutter will begin on 23 April 2007. The LRIS blue channel shutter should be available in time for the next run which starts 7 May 2007.

For now, the trapdoor may be used as a shutter for the blue side.In this mode, red side data may still be acquired simultaneously. Your SA will review the scripts with you upon your arrival, but please see trapdoor_as_shutter.html for a brief description of the observing scripts.

Observers who used LRIS during April 2007 should carefully examine blue side data to ensure the shutter was not stuck open during their observations.

LRIS Red channel missing columns (2007-April-13)

LRIS Troughput Measurements (2007-Mar-12)
To regularly monitor the throughput of LRIS, observers will be asked for permission to take spectrophotometric observations of a designated standard star during twilight. A script called observe_flux_ standard: verifies that the telescope is pointed to a known LRIS flux standard star by checking DCS keyword TARGNAME; checks the current grating and build appropriate exposure list; saves current LRIS configuration; checks that SKYPA is correct for this star; sets the correct pointing origin (SLITB); configures for slitless spectroscopy; acquires exposures; and restores the observer's configuration. Before running the script the observer should be set up to acquire spectroscopic measurements on the red side (blue side optional) and the OA should point to one of the standards in the target list called 0000_throughput_standards.

ADC status (2007-March-7)
For a description of the current status on the ADC please see: ADC Current Status

LRIS Blue side shutter replaced (2006-Sept-7)
The blue channel dewar is installed in LRIS following the replacement of the blue channel shutter. The shutter that was replaced failed due to ware between a pivot pin and arm used to open and close the shutter. The new shutter was tested and is operating normally.

As a result of the shutter replacement, the blue channel dewar was realigned. As a result of the realignment, the FOV is shifted slightly along rows and columns. This shift along the horizontal axis for LRIS is roughly 70 pix. The shift is benefits LRIS observers because it moves the bar closer to the center of the gap between the blue channel ccds. Observer's may notice that slits fell in the gap during May to Aug because the dewar was slightly misaligned.

The dewar is rotated 0.1-0.3 degrees such that the edges of the imaging FOV are not parallel to the horizontal and vertical array axis.

Grating station 3 is still unavailable. Observers will be limited to three grating choices until the grating station three is fixed. Thus, the fourth grating listed in an observer configuration form will not be installed as it is assumed to have the lowest priority.

LRIS Blue side shutter is stuck open (2006-Aug-29)
The LRIS blue channel shutter has failed again, and work to replace the shutter began on 28 Aug 2006. The LRIS blue channel should be available in time for the next run which starts 15 Sept. 2006. The shutter that recently failed was a rebuilt shutter. We are replacing it with a new shutter. We are replacing the shutter this week, and the blue channel dewar will be re-aligned next week.

Observer who used LRIS during 21-27 Aug. 2007 should carefully examine blue channel data to ensure the shutter was not stuck open during their observations.

Blue channel FOV is slightly misaligned (2006-Aug-18)
The dichroics used to redirect light toward the blue channel are slightly misaligned. The misalignment results in all slits being shifted roughly six arcseconds to the right as seen on the CCD mosaic. This means that some portion of a slit very near the slitmask bar will fall in the gap between the two Blue channel CCDs. If this shift significantly disrupts the primary science with your slitmasks for the run starting 21 August, you may redesign and submit a slitmask for milling.

Grating station three Fixed (2006-Aug-10)
Grating station 3 is back in action. The encoder used to help define the position of the grating tilt failed. The failure mode was that the encoder could not find home. The encoder was replaced with a spare. Observers should find that the wavelength offsets for grating station three that were measured during past runs may no longer apply. Observers may again select up to four different gratings for their run.

Grating station three failed (2006-Jul-18)
Grating station 3 is temporarily out of service. The grating station failed 18 July, and has been taken out of service. Observing teams are limitted to 3 gratings until the grating station is fixed.

Dewar Service and Focus Change (2006-Jul-18)
The LRIS dewar was removed from the instrument so that summit staff could install a new dewar window, replacing the original which was broken during the shutter repair in May. The new window appears to work as desired. Following the work, we've been adjusting the "tilt" position of the LRIS blue dewar to improve the uniformity of focus across the field of view. Currently the unformity in the blue CCD Y direction (along CCD columns) is good, but the CCD X direction (along rows) shows a gradient of about 100 microns from side to side. We judge this to be an insignificant amount of defocus. Observers should be aware that the nominal blue-side focus value has changed significantly and is now approximately -2450. This should be used as the central focus value when acquiring focus images with the focus loop script.

New Web Pages Released (2006-May-15)
There is a new look and feel to the LRIS instrument web pages. The new web pages include information on the LRIS blue channel that was previously lacking with the old pages. Procdures were updated, conflicts in the information or procedures were resolved, and the page information was re-organized. The LRIS documatation is grouped into three primary stages: Pre-Observing when you will require information for preparing proposals and materials in advance of an observing run; Observing when you will be at the telescope, and Post-Observing when you will need to process and analyze data in preparation for publishing.
The old web documenttion is obsolete and is no longer available. Please report any errors, missing information, clarification requests, and any problems with the pages to the lris instrument master.

LRIS Blue side shutter is stuck open! (2006-Apr-25)
The LRIS blue side shutter has failed and work to replace the shutter will begin on 5 May 2006. The shutter is stuck open which permits observing. Several new scripts are available to observers for use during afternoon setup and science acquisition. The trapdoor is currently being used as a shutter for the blue side. Red side data may still be acquired simultaneously. Your SA will review the scripts with you upon your arrival, but please see trapdoor_as_shutter.html for a brief description of the observing scripts.

LRIS Configuration Form Updated (2005-Nov-16)
A new version of the LRIS configuration form is available at . The new form is customized to the number of nights in your observing run, links to the slitmask database, and populates dropdown lists with facility and user masks. For LRIS, the blue side elements are listed, and a default configuration is presented. When prepairing for upcoming LRIS and DEIMOS runs, please use the new configuration forms.

IRAF script updates (2005-May-19)
A new version of is available in the keck.lris package to handle the 4-amp case (old version only worked with 1 or 2 amps). The script was tested with both red and blue images. Please report any problems with the script to the LRIS master. Chris Fassnacht is the author of the new version of

Even Older News ...